OpenStudy (anonymous):

let R^4 have the Euclidean inner product. Find 2 unit vectors that are orthogonal to all three of the vectors u=(2,1,-4,0), v=(-1,-1,2,2)and w =(3,2,5,4) . Guide me, please

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

what you tried?

OpenStudy (anonymous):

try to find a vector which satisfy the requirement

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I would use Gram-Schmidt

OpenStudy (anonymous):

common, I tell you I am not allowed to use it now, just after tomorrow class, i will. but not now.

OpenStudy (anonymous):

anyway by commun sence, you would like a vector (a,b,c,d) which has no component in u,v,w directions. So use inner product to find it's projection on each of the u,v,w vectors and rest it from this vector

OpenStudy (anonymous):

ok, thanks, let me try and tell you the result. if it fix your way by applying Gram. just say yes to mine.

OpenStudy (anonymous):

let \[x=\left[\begin{matrix}a\\b\\c\\d\end{matrix}\right]\\ u\cdot x=0,\quad v\cdot x=0\quad w\cdot x=0\quad x\cdot x=1 \] the four oprthognoality conditions will give you with four simultaneous equations.

OpenStudy (anonymous):

i think so, too.

OpenStudy (anonymous):

why did I get the medal? heres to myko.

OpenStudy (anonymous):

you rather help other than me. but at the end up, you are here. forgive you